### v1

# PromptHSI: Universal Hyperspectral Image Restoration Framework for Composite Degradation

Abstract
Recent developments in All-in-One (AiO) RGB image restoration and prompt learning have enabled the representation of distinct degradations through prompts, allowing degraded images to be effectively addressed by a single restoration model. However, this paradigm faces significant challenges when transferring to hyperspectral image (HSI) restoration tasks due to: 1) the domain gap between RGB and HSI features and difference on their structures, 2) information loss in visual prompts under severe composite degradations, and 3) difficulties in capturing HSI-specific degradation representations through text prompts. To address these challenges, we propose PromptHSI, the first universal AiO HSI restoration framework. By leveraging the frequency-aware feature modulation based on characteristics of HSI degradations, we decompose text prompts into intensity and bias controllers to effectively guide the restoration process while avoiding domain gaps. Our unified architecture excels at both fine-grained recovery and global information restoration tasks. Experimental results demonstrate superior performance under various degradation combinations, indicating great potential for practical remote sensing applications. The source code will be publicly released.
Corresponding author, email: cchsu@gs.ncku.edu.tw
1 Introduction
Figure 1: The All-in-One and universal HSI restoration framework with text-prompt guidance, named PromptHSI, for composite degradation Specifically, informative textual features extracted by pre-trained vision-language model are fed into adapter and then decomposed into an intensity and bias embeddings for degradation-specific restoration with effective cascade guidance.
Hyperspectral images encompass a greater number of spectral bands than RGB images, offering extensive spectral details [44]. Because materials reflect different spectral bands, each substance can have a unique signature for material identification [1, 37, 44, 32]. HSI can capture these valuable information and has broad application value and potential. However, because of the limitation of the device and bandwidth, observed HSIs are usually damaged by different types of degradation from camera sensing and transmission. Thus, HSI restoration is a crucial pre-processing step to recover degraded pixels from flawed observations.
Previous works tended to be based on HSI structures, low-rank sparsity prior [43, 11, 42] for specific degradations such as Gaussian noise [25, 49], stripe effect [24, 22], spatial blurring [33, 19], and cloud occlusion [35, 48]. However, in real-world scenario, HSIs are easily prone to multiple types of degradation, causing these methods to be limited. When HSIs are subject to multiple degradation, some methods based on deep unfolding network [4, 34] or optimization [29, 28, 52, 42] perform significantly below expectations and lack flexibility to use for real-world application.
With the development of AiO RGB image restoration methods [39, 5, 31, 10, 13], which can address multiple degradations in a single network, there is a potential to resolve this situation. However, due to the significant domain gap between the HSI and RGB images, these AiO methods cannot be transferred to HSI restoration. In particular, the visual-prompt approaches [39, 5] are influenced by severe and diverse corruptions. Furthermore, while text-prompts [10, 13] have emerged as an effective approach for controlling All-in-One restoration models, the textual feature representations encoded by widely-used vision-language models (VLMs) pre-trained on RGB-text datasets, such as CLIP [40], exhibit significant domain gaps with HSI features. Given these limitations, leveraging prompt learning frameworks for HSI processing remains an open challenge.
Considered as a whole, the essential questions are Can we leverage prompt learning and vision-language model (VLM) frameworks to address critical challenges in hyperspectral image processing for practical applications, such as achieving composite degradation restoration or controllable restoration through a single unified model? The response is affirmative.
To meet these challenges, we propose PromptHSI, a frequency-aware and prompt-guided framework tailored for HSI’s unique spectral structure and degradation types. Unlike AiO RGB image restoration methods, it ensures spectral fidelity and adaptive degradation response through novel feature modulation strategies. Our universal restoration architecture effectively handles both fine-grained recovery (e.g. spatial-spectral super-resolution) and global information restoration (e.g. decloud, inpainting) within a single model; a frequency-aware feature modulation strategy that decomposes text-prompts into intensity and bias embeddings based on frequency-domain analysis, dispelling the domain gap between VLM and HSI features. In summary, our main contributions can be listed as follows:
  • We present the first universal HSI restoration framework capable of handling composite degradations. Unlike existing HSI restoration methods that suffer from performance drops in composite degradations, our approach maintains consistent high performance across various degradation combinations.
  • We propose a novel text-to-feature modulation mechanism that effectively leverages frequency-domain characteristics of HSI degradations. By decomposing textual features into intensity and bias controller embeddings, our method achieves superior control over the restoration process, outperforming state-of-the-art AiO RGB image restoration methods on composite degradation.
  • Extensive experiments demonstrate that our method not only excels at HSI restoration under various degradation combinations but also exhibits strong generalization to unseen degradation types through flexible text-prompt control. The framework’s effectiveness is thoroughly validated through comprehensive ablation studies and comparisons with existing approaches.
Figure 2: Illustration of the degraded HSI and corresponding description synthesis pipeline. Each degradation is activated with probability $p$. Note that we set $p$ = $0.5$ for all experiments in this paper. To ensure diversity and realistic in the synthetic dataset, each activated degradation selects and applies one of its sub-degradations to the inputted HSI. (More detailed information and setting can be found in Supplementary.)
2 Related Works
Figure 3: Left, displayed from top to bottom, are HSIs with specific degradation alongside the Fourier spectra of the residual images, which are derived from subtracting the degraded images from the ground truth. Right, the diagram presents the mean values of the Fourier spectra based on the square length indicated on the x-axis. It is evident that varying degradations focus on distinct frequency subbands.
2.1 Hyperspectral Image Restoration
Conventional approaches are based on the structural properties of HSI, such as low-rank and sparse priors [43, 11]. Optimization-based methods [52, 36, 42, 6] were initially used for HSI restoration. With the progress of deep learning, HSI can be effectively recovered by imposing appropriate constraints and suitable design for task-specific restoration. Unlike RGB image restoration, HSI restoration considers not only global information aggregation but also spatial-spectral correlation at locality. Several works against different types of restoration model, such as denoising [24, 38, 15] for Gaussian noise and impulse noise, inpainting [6, 52, 29] for stripe effect, and spatial blurring [20, 26, 18] or limited spectral resolution [14, 2, 3].
However, due to multiple or composite degradations that jointly affect structural information and fine-grained details of a given HSI, restoration methods are easily compromised, consistently leading to suboptimal performance.
2.2 All-in-One RGB Image Restoration
A recently emerging paradigm in RGB image restoration involves designing single models to restore different types of degradation [12, 7, 23, 39, 10, 5, 13]. Current techniques, as a type of multi-task learning, utilize various frameworks to manage inputs and outputs for multiple tasks with prompting methods, facilitating effective information sharing between them and corresponding degradation. Specifically, AirNet [23], the first proposed AiO image restoration method, employed contrastive learning to train the encoder to learn degradation representations and restoration with a single network. PromptIR [39] introduced the visual prompt for the degradation perception and information integration within the decoding process without a priori needs. InstructIR [10] followed the text-prompt as instruction to guide the network, archiving great results with controllability. HAIR [5] is based on the hypernetwork, which aims to generate network parameters for corresponding degradations. Recently, OneRestore [13] has extended AiO restoration from multiple degradation to composite degradation, which can address the various corruption even they exists in a same image simultaneously.
Although these methods are effective, these solutions are limited in HSIs. Due to domain gap and composite degradation, it is difficult to effectively learn the compact representation of degradation and perform effective restoration.
3 Preliminary and Motivation
3.1 Unified HSI Restoration Network Design
For comprehensive restoration using a single network, employing an architecture suitable for various degradations is crucial [8, 47]. In RGB image restoration, two classic architectures predominate: ❶ U-shape encoder-decoder [46, 45] and ❷ plain residual-in-residual [27, 17, 50]. The former excels at capturing general representations and information, being widely used in denoising, deblurring, and inpainting, but pays less attention to fine details or repeat patterns. The latter specializes in capturing detailed spatial features at the fine-grained level, serving as the backbone for super-resolution tasks [9, 41, 51].
Previous methods have overlooked the compatibility between these two restoration models, and current AiO restoration approaches [39, 5, 13] typically focus on the former rather than integrate the joint strength of both types of network design, limiting their versatility. Besides, this limitation also extends to HSI restoration models [33, 19, 21]. Therefore, we integrate a plain residual-in-residual block into the U-net cascade decoder in the PromptHSI to maintain the advantages of both designs without compromising. The versatility of the network can be ensured accordingly.
3.2 Degradations Analysis in Frequency Domain
Unlike RGB image restoration, HSI restoration requires careful consideration of spatial-spectral correlations across hundreds of bands. Our key insight comes from analyzing how different types of degradation manifest in the frequency-domain through Fourier analysis. Given a clean HSI $I_{\text{clean}}$, its Fourier transform can be expressed as:
$F(I_{\text{clean}})=F_{s}(I_{\text{clean}})+iF_{c}(I_{\text{clean}})$ (1)
where $F_{s}$ and $F_{c}$ represent the sine and cosine components respectively. Through empirical analysis, as shown in Figure 3, we observed that different types of degradation exhibit distinct patterns in how they affect these frequency components.
Specifically, the relationship between degraded HSI $I_{\text{degraded}}$ and clean HSI $I_{\text{clean}}$ in frequency-domain can be characterized by an affine transformation:
$F(I_{\text{degraded}})=\hat{\lambda}\odot F(I_{\text{clean}})+\hat{\mu}=(1+\lambda)\odot F(I_{\text{clean}})+\mu$ (2)
where $\odot$ denotes element-wise multiplication, $\lambda$ and $\mu$ are degradation-specific parameters that: $\lambda$ represents the scaling factor that modulates frequency amplitudes, $\mu$ is the bias term that shifts the frequency distribution.
Different types of degradation show characteristic patterns in $\lambda$ and $\mu$, such as ❶ Blur mainly affects high frequencies: $\lambda_{high}$ $\ll$ $\mu_{low}$. ❷ Noise adds uniform perturbations: $\mu_{noise}$approx constant $c$. ❸ Missing information creates structured patterns in both $\lambda$ and $\mu$.
4 Proposed Method
Figure 4: The architecture of the proposed PromptHSI, which is designed to tackle the composite degradations within a single universal model. First, we design a network based on U-net-like encoder-decoder architecture [46, 45] for capturing global spatial-spectral information, while combining Residual-Dense Group (RDG) and Fast-Residual-Dense Block (FRDB) presented in DRCT [17] and RTCS [16], respectively, in the proposed PGFAB for spatial-spectral enhancement. Afterwards, text-prompt-guidance and frequency-aware feature map modulation is proposed to jointly better capture the representation of composite degradation with controllability.
Overview. Figure 4 illustrates the overall architecture of PromptHSI, which consists of three key components designed to address the challenges in universal HSI restoration: ❶ A unified U-shape network which contains Spatial-Spectral Aware Feature Extraction Block (SSAEFB) and Propmpt-guided Feature Aggregation Block (PGFAB) that effectively captures both enhanced spatial-spectral fine-grained details and global contextual information from degraded HSI; ❷ A Prompt-Guided Feature Modulation (PGFM) that effectively bridging the domain gap between textual features and HSI features through frequency-domain modulation, while avoiding direct cross-domain feature alignment and information loss.
4.1 Hyperspectral Image Restoration Pipeline
Our restoration pipeline consists of two key components: a SSAEFB for efficient feature extraction, and a PGFAB for controllable restoration with text-prompt guidance.
The encoder exploits the low-rank and sparse prior of HSIs through depth-wise separable convolution, effectively reducing computational complexity while preserving spectral characteristics. PGFAB then utilizes dual branches for spatial and spectral enhancement, with cross-attention mechanisms facilitating comprehensive feature integration.
Spatial-spectral-aware encoder. In conventional RGB restoration and other vision tasks, standard $3\times 3$ convolutions are widely considered to be sufficiently versatile and effective feature encoders. However, considering the low-rank characteristics of HSI, where substantial redundancy exists between different spectral bands, utilizing standard convolutions may reference unnecessary information from distant spectral bands, leading to suboptimal performance. Therefore, encoders based on lightweight convolutions prove sufficiently effective while further reducing parameter size to meet practical application requirements.
Specifically, for the input HSI $I_{\text{degraded}}$, we first interpolate the spectrally discontinuous regions (caused by the stripe effect) to facilitate the learning process. Subsequently, we apply standard convolution layers to obtain shallow feature maps.
Afterwards, SSAFEBs is based on a depth-wise separable convolution to extract features while inherently taking advantage of this low-rank prior, thereby reducing computational complexity without compromising performance [16].
Propmpt-guided feature aggregation block. The features captured by SSAFEB are combined with the informative text-prompt guidance from PGFM (which will be discussed in the next subsection). Using effective guidance, the network is allowed to dynamically activate the corresponding weight for HSI restoration according to different types of degradation. To capture detailed information, PGFAB utilizes two distinct branches: one focuses on spatial dimension and the other on spectra. These branches are eventually integrated for feature aggregation.
Spatial and spectral feature enhancement. The output of PGFM is first improved by consecutive RDG [17] and FRDB [16] to improve fine-grained details. Then these features are interacted by cross-attention layer, which aims to comprehensively integrate the spatial-spectral feature information from different branches across all dimensions. Specifically, it can be expressed as follows:
$\{Q_{\alpha},K_{\alpha},V_{\alpha}\}=\mathcal{F}^{\alpha}_{qkv}(F_{\alpha}),\quad\{Q_{\beta},K_{\beta},V_{\beta}\}=\mathcal{F}^{\beta}_{qkv}(F_{\beta}),$ (3)
where $F_{\alpha},F_{\beta}$ denote features from the spatial and spectral enhancement block. Subsequently, we exchange the queries $Q$ of two branches for spatial-spectral interaction:
$F_{f}^{\beta}=softmax(\frac{Q_{\alpha}K_{\beta}}{d_{k}})V_{\beta},\quad F_{f}^{\alpha}=softmax(\frac{Q_{\beta}K_{\alpha}}{d_{k}})V_{\alpha},$ (4)
where $d_{k}$ is the scaling factor. Subsequently, we concatenate the results obtained by the cross-attention through $F_{f}^{\gamma}=Concat({F_{f}^{\beta},F_{f}^{\alpha}})$ to obtain the integrated features.
$\{Q_{\gamma},K_{\gamma},V_{\gamma}\}=\mathcal{F}^{\gamma}_{qkv}(F_{f}^{\gamma}),\quad\hat{F_{f}^{\gamma}}=softmax(\frac{Q_{\gamma}K_{\gamma}}{d_{k}})V_{\gamma},$ (5)
Finally, we employ spectral-spatial self-attention for feature co-enhancement to improve representation capability.
Through a dual-branch architecture, we integrate text-prompt guidance effectively across both spatial and spectral dimensions. By incorporating plain-residual-in-residual [27, 17, 16] modules within the U-shape cascade decoder [46, 8], we enhance fine-grained image restoration while integrating global information, thereby effectively addressing various degradations for HSI restoration.
Table 1: Overall performance evaluation for composite degradation and complexity comparison. To simulate composite degradations, we maintain a fixed $p$ of 0.5 for triggering individual degradation types, as demonstrated in Figure 2, for all experiments. Note that Cr. means Controllable or Not. As for the complexity parts, M and G indicate $10^{6}$ and $10^{9}$, respectively. The best, second-best and third-best results are highlighted in bold-red, bold-blue, and bold-green, respectively.
[…]
Input - - - - $\times$ 15.2454 17.7463 0.0643 33.7220 - -
HSIDwRD [49] ICCV 2021 Denoising HSI $\times$ 20.6270 12.7891 0.0364 97.3601 23.62M 5054.16G
SERT [25] CVPR 2023 Denoising HSI $\times$ 16.1657 15.9853 0.0348 29.2053 2.74M 176.06 G
SRDNet [33] TGRS 2024 Super-Res. HSI $\times$ 19.0520 11.9590 0.0320 16.5861 1.86M 3771.80G
SNLSR [19] TIP 2024 Super-Res. HSI $\times$ 17.8785 20.4389 0.0445 19.1267 1.75M 14.20G
One-to-One SGNet [35] ISPRS 2022 Decloud HSI $\times$ 18.0694 12.9476 0.0359 17.9773 0.54M 57.39G
AACNet [48] TGRS 2023 Decloud HSI $\times$ 17.6279 14.5962 0.0324 26.2115 2.76M 171.23G
SST [24] AAAI 2023 Inpainting HSI $\times$ 18.9792 10.9860 0.0244 26.9934 4.30M 281.72G
HSDT [22] ICCV 2023 Inpainting HSI $\times$ 18.0665 11.3451 0.0372 68.4451 0.51M 67.55G
AirNet [23] CVPR 2022 - RGB $\times$ 23.6647 7.5970 0.0233 16.4194 6.07M 245.86G
All-in-One PromptIR [39] NeurIPS 2023 - RGB $\times$ 17.2667 9.6437 0.0405 29.2768 33.18M 132.07G
for specific HAIR [5] arXiv 2024 - RGB $\times$ 20.5779 14.4329 0.0275 24.4211 7.86M 42.51G
InstructIR [10] ECCV 2024 - RGB $\checkmark$ 20.3106 14.5727 0.0306 27.5454 16.02M 17.28G
All-in-One OneRestore [13] ECCV 2024 - RGB $\checkmark$ 17.3321 15.0320 0.0271 52.1162 5.85M 12.05G
for composite PromptHSI (Ours) - - HSI $\checkmark$ 26.4432 6.0601 0.0185 8.3566 26.14M 135.83G
Figure 5: Visualization results for HSI restoration methods. top: the reconstructed HSI, bottom: the residual image derived from Ground-truth and reconstructed result. Most of state-of-the-arts HSI restoration methods fail to deal with composite degradations.
4.2 Prompt-Guided Feature Modulation
Directly using text features from VLMs for HSI restoration faces several critical challenges. Firstly, VLM features are trained on RGB-text pairs, lacking understanding of spectral characteristics and capturing complex spectral degradations. Secondly, their statistical distribution are mismatched because VLM features follow RGB image statistics, but HSI has essentially different data distributions, and direct adaptation requires complex domain translation.
Instead of directly bridging the HSI-text domain gap, our proposed PGFM addresses the domain gap by decomposing the problem into two easier sub-tasks. ❶ frequency-pattern matching with task descriptor and ❷ distribution adjustment by novel frequency-aware feature modulation mechanism. This strategy aim to selectively enhance spectral fidelity across frequency bands, thereby overcoming the limitations of text-guidance for avoiding direct cross-domain alignment between text features extracted from VLMs pre-trained on RGB-text domains and HSI.
Task descriptor generation. Given a text prompt $T_{text}$ that provides the corresponding degradation description to guide the HSI restoration network to obtain the specified restoration result. In this study, we employ CLIP [40], which integrates vision and language information and performs effectively in text feature extraction. We freeze the pre-trained CLIP to maintain great linguistic consistency and add additional MLP layer $\Phi_{t}^{adpt.}$ as the feature adapter for pattern matching. With {}$\{\cdot\}_{e}$ denoting the frozen weights, this process can be expressed as:
$\footnotesize F_{text}=\Phi_{t}^{adpt.}(\{\mathcal{F}^{CLIP}_{text}\}_{e}(T_{text})),$ (6)
where $F_{text}$ denotes the task descriptors. With the task descriptors available, the next challenge is how to effectively incorporate this guidance into the HSI restoration process.
Frequency-aware feature modulation. Based on the frequency domain analysis, as illustrated in Section 3.2, we design our feature modulation mechanism to mimic how degradations affect different frequency components. Given an input feature map $F_{f}$, we first decompose it into frequency components:
$F_{f}=F_{f}^{l}+F_{f}^{h},$ (7)
where $F_{f}^{l}$ and $F_{f}^{h}$ represent low and high frequency components, obtained through frequency-domain filtering.
Afterward, we design the MLP layer $\Psi_{m}^{i}$ to extract this connection with the intensity and bias embeddings for feature modulation. Therefore, it can be obtained:
$\lambda_{m}=\Psi_{m}^{I}(F_{text}),\ \mu_{m}=\Psi_{m}^{II}(F_{text}),$ (8)
where $\Psi_{m}^{I}$ and $\Psi_{m}^{II}$ are the chunk operations of $\Psi_{m}$ to form the intensity and bias controller embeddings. To address the spectral consistency challenge in HSIs, our frequency-aware feature modulation adopts an affine transformation model tailored for spectral data, formulated as:
$\hat{F}_{f}=(1+\lambda_{m}^{l})\odot F_{f}^{l}+(1+\lambda_{m}^{h})\odot F_{f}^{h}+\mu,$ (9)
where Intensity controller $\lambda$in [$\lambda_{m}^{l}$, $\lambda_{m}^{h}$] modulates the amplitude of high-frequency components to address noise, while Bias controller $\mu$ shifts the low-frequency baseline to compensate for blurring and band-missing effects. This design ensures frequency-specific adjustments that RGB AiO frameworks lack.
To validate the efficacy of this mechanism, we performed ablation studies demonstrating improved spectral fidelity and PSNR consistency across spectral bands, underscoring the need for spectral domain modulation to manage complex HSI degradations effectively.
4.3 Loss Function
Given the wide range of spectra inherent in HSI, we utilize only the $\ell_{1}$ loss for training. However, this strategy may not adequately capture the variability across different bands. Consequently, the $\ell_{1}$ loss might inadvertently focus on minimizing bands with larger values, affecting the spectral integrity of specific bands.
Spectral angle mapper loss. SAM metric is also critical for material identification and less sensitive to intensity variations. The $\ell_{\text{SAM}}$ loss prioritizes angular discrepancies over magnitudes, offering a reliable metric to evaluate the spectral fidelity of the reconstructed HSI.
$\ell_{\text{SAM}}=\frac{1}{N}\sum_{n=1}^{N}\cos^{-1}\left(\frac{I_{n}^{T}I^{*}_{n}+\epsilon}{\|I_{n}\|_{2}\cdot\|I_{n}^{*}\|_{2}+\epsilon}\right),$ (10)
where $I_{n}$ and $I^{*}_{n}$ represent the $n$-th spectral vector of the ground-truth and the reconstructed HSI, respectively. The term $\epsilon$ is included to avoid divergence.
Stationary wavelet transform loss. Wavelet sub-band alignment is able to manage various aspects of information encoded in HSIs. On the other words, it facilitates more detailed supervision in the frequency-domain.
$\ell_{\text{SWT}}=\sum_{j}\lambda_{j}\big{\|}\text{SWT}(I)_{j}-\text{SWT}(I^{*})_{j}\big{\|}_{1},$ (11)
where $\lambda_{j}$ represents the count of sub-bands decomposed via stationary wavelet transformation.
Band-wise MSE loss. The band-wise disparity helps stabilize the network optimization process and is advantageous to achieve convergence, balancing restoration quality across all spectra. The Band-wise MSE loss can be simply defined as follows:
$\ell_{\text{BMSE}}=\frac{1}{N}\sum_{n=1}^{N}{\|I_{n}-I_{n}^{*}\|_{2}},$ (12)
Total loss. Ultimately, the total loss function for training the proposed PromptHSI $\ell_{\text{total}}$ can be expressed as:
$\ell_{\text{total}}=\lambda_{1}\ell_{\text{1}}+\lambda_{2}\ell_{\text{SAM}}+\lambda_{3}\ell_{\text{SWT}}+\lambda_{4}\ell_{\text{BMSE}},$ (13)
where $\lambda_{1,2,3,4}$ serve as the weighting coefficients to equilibrate $\ell_{1}$, $\ell_{\text{SAM}}$, $\ell_{\text{SWT}}$, and $\ell_{\text{BMSE}}$, set at 1, 0.1, 0.01, 0.01 throughout all experiments, respectively.
5 Experimental Results
Figure 6: Comparison for AiO RGB image restoration methods. For the text-prompt-based methods [10, 13] and our PromptHSI, we employ the standard text-prompts presented in this study for training network and inference for related experiments.
Figure 7: Comparison of t-SNE visualization results for the degradation embeddings of the proposed PromptHSI trained on single degradation. (a) for multiple degradation, (b) for composite degradatation. The proposed PromptHSI have demonstrated its discriminative capability in identifying different degradation even unseen degradation and its combination during testing-time.
Table 2: Ablation of integration strategy.
Strategy PSNR / SAM / RMSE / EGRAS
Intensity-only 26.0070 / 6.3636 / 0.0187 / 9.7608
Bias-only 25.6384 / 6.3963 / 0.0190 / 12.4992
Weighted-sum 25.7284 / 6.3561 / 0.0188 / 14.3605
Intensity & Bias 26.4432 / 6.0601 / 0.0185 / 8.3566
Table 3: Effect of prompting method.
Prompt type Cr. PSNR / SAM / RMSE / EGRAS
RGB to VLM $\times$ 24.1273 / 7.6236 / 0.0257 / 14.5208
Hybrid to VLM $\checkmark$ 25.2385 / 6.8974 / 0.0191 / 11.4359
PGB & PIM [39] $\times$ 23.0567 / 9.4862 / 0.0287 / 24.5575
Text to VLM $\checkmark$ 26.4432 / 6.0601 / 0.0185 / 8.3566
Table 4: Ablation of network component.
Without PSNR / SAM / RMSE / EGRAS
Full model 26.4432 / 6.0601 / 0.0185 / 8.3566
PGFM 25.4905 / 6.6348 / 0.0205 / 13.2072
adapter 25.8686 / 6.3561 / 0.0188 / 14.3605
enhancement 23.7511 / 7.9918 / 0.0262 / 16.0007
Experiment Setup. This study employed the AVIRIS dataset provided by Lin et al. [30, 18], an extensive HSI dataset, with numerous bands spanning diverse geographical areas for experiments. It includes 2,078 HSIs, which were randomly divided into training, validation, and testing sets. The training set consisted of 1,678 images and both the validation and testing sets contained 200 images each. As illustrated in Supplementary, all clean HSI was pre-processed to synthesize the composite degraded HSIs and corresponding text descriptions.
5.1 Quantitative Results and Analysis
To evaluate the performance of PromptHSI, we compared it with eight other supervised HSI restoration methods for specific degradation, including HSIDwRD [49], SERT [25], SRDNet [33], SNLSR [19], SGNet [35], AACNet [48], SST [24], and HSDT [22] in the synthesized AVIRIS dataset with composite degradations. Because there is no AiO HSI restoration method, we compare four AiO RGB image restoration methods in the experiments: AirNet [23], PromptIR [39], HAIR [5], InstructIR [10], and OneRestore [13]. The model performance was measured using the four metrics, including PSNR, SAM, RMSE, and ERGAS. Table 1 presents the quantitative results.
Ours vs. Task-specific HSI restoration methods. All peer-methods fail to effectively recover the degraded HSIs due to their ability to learn different or unseen degradations. In contrast, PromptHSI can restore better even without text-prompt, as shown in Table 4, which exhibits the capability of our unified restoration network design.
Ours vs. All-in-One RGB restoration methods. Unlike AiO RGB image restoration methods that rely on uniform feature modulation, our PromptHSI framework’s frequency-aware modulation dynamically tailors restoration to each spectral band, thereby achieving cross-band consistency crucial for HSI applications.
In general, our PromptHSI exceeds other AiO methods. Firstly, most of existing methods struggle to handle composite degradations. As discussed previously, both data and features extracted by pre-trained CLIP [40] exhibit domain gaps with HSI. Therefore, these models are not effective in transferring to AiO HSI restoration.
Specifically, OneRestore [13] incorporated task description features into its encoder block, which hindered its capacity to accurately capture the spatial-spectral correlation in HSI because of the domain gap with HSI features, resulting in decoding failures. Moreover, it lacks considerations for tasks requiring high-frequency detail reconstruction, potentially limiting its effectiveness in the spatial or spectral super-resolution task.
5.2 Ablation Studies and Analysis
Figure 8: Visualization results for the controllability. Top, for single-specific degradation, the plots in the first row represent the degraded HSI, while the second row shows residual maps, obtained from the reconstructed HSI, with or without the guidance provided by the text-prompt. Bottom, in the case of composite degradation, the top row of plots depicts the recovered HSI, while the residual maps are obtained from the reconstructed HSI using the specified text-prompt guidance with the ground-truth.
The aforementioned experimental results demonstrate the effectiveness of our proposed PromptHSI framework under composite degradations. We further conduct ablation studies to investigate the impact of different components and strategies.
Integration strategy of feature-map modulation. As shown in Table 4, our proposed frequency-aware feature modulation demonstrates superior effectiveness compared to the weighted-sum manner, where we implemented an MLP layer for dynamic weighted averaging of inputs and outputs. The intensity-only configuration outperforms the bias-only, with a notably larger advantage in the ERGAS metric, suggesting that this integration strategy effectively guides network restoration across most scenarios.
Visual-prompt or text-prompt. The type of prompting is crucial, as it determines its ability to recognize various degradations and effectively guide and control model behavior. In the first configuration, we select RGB bands from HSIs as visual prompts, encoding them through CLIP [40] and using the proposed PGFM to guide the model. Secondly, we encode both RGB bands and textual prompts to generate hybrid task descriptors for guidance. In the final configuration, we remove PGFM and use the PIM and PGM block proposed in PromptIR [39] for the generation of visual prompts and interaction of features.
The experimental results are presented in Table 4. Due to multiple or composite degradations in the spatial-domain leading to serious information loss, it became difficult to obtain compact task descriptors or valid visual prompts, ultimately resulting in diminished prompt effectiveness and compromised restoration. The use of textual features helps to circumvent these issues and achieves robust prompting guidance through semantic information.
Ablation of network component. To understand the impact of the network components, we performed ablation studies by removing PGFM, adapter and spatial spectral enhancement from the network, with results shown in Table 4. Even without textual features, our proposed PromptHSI outperforms other peer methods in handling composite degradation, demonstrating the effectiveness and universality of our network architecture. The adapter proves crucial in bridging the domain gap between the VLM and HSI features via pattern matching, leading to improved global reconstruction performance, particularly evident in the ERGAS metric. Lastly, we can observe that there is a drop with or without spatial-spectral enhancement, which represents that feature enhancement is essential for details recovery within cascade decoder.
These results demonstrate PromptHSI’s superior capability in maintaining spectral fidelity across multiple degradation types, highlighting the necessity of frequency-aware modulation, particularly when compared to AiO RGB methods that exhibit substantial performance degradation in the HSI domain. In summary, PromptHSI offers a universal and controllable solution for addressing composite degradation in real-world applications.
6 Conclusion
In this paper, we present PromptHSI, the first universal framework for hyperspectral image restoration that effectively addresses composite degradations, which conventional HSI restoration methods have failed to handle. Unlike existing All-in-One RGB image restoration approaches that struggle with domain gaps between RGB and HSI features and structures, our method is specifically designed to address HSI-specific attributes and commonly encountered degradation patterns. Then, our method is controllable through text-prompt guidance, based on a novel frequency-aware modulation strategy that effectively bridges the gap between text and HSI through frequency pattern matching and distribution adjustment in the frequency-domain. Our approach demonstrates superior performance across various types of degradation and their combinations, significantly outperforming existing methods.